A distribution of values is normal with a mean of 33 and a standard deviation of 91.4.
Find P26, which is the score separating the bottom 26% from the top 74%.
P26 =
Enter your answer as a number accurate to 4 decimal places.

Answer by Boreal(15235)

want z-value for 26th percentile
this is -0.6433
z=(x-mean)/sd
-0.6433=(x-33)/91.4
-58.8018=(x-33)
x=-25.8018